1. For what values of h the vectors



⟶ [ 1 ⟶ [ -1 ⟶ [ 5

u1 = -3 u2 = 9 u3 = -7

-2 ] -6 ] h ]


are linearly independent? (Show all working)


1
Expert's answer
2022-03-21T02:00:32-0400

"\vec{a}=(1,-1,5), \vec{b}=(-3,9,-7), \vec{c}=(-2,-6,h)\\
\lambda_1\vec{a}+\lambda_2\vec{b}+\lambda_3\vec{c}=\vec{0}\\
\lambda_1=\lambda_2=\lambda_3=0\\
\lambda_1(1,-1,5)+\lambda_2(-3,9,-7)+\lambda_3(-2,-6,h)=\\=(0,0,0)\\
\lambda_1-3\lambda_2-2\lambda_3=0\\
-\lambda_1+9\lambda_2-6\lambda_3=0\\
5\lambda_1-7\lambda_2+h\lambda_3=0\\
\Delta=
\begin{vmatrix}
 1& -3&-2 \\
 -1 & 9&-6\\
5&-7&h
\end{vmatrix}=9h+90-14+\\+90-3h-42=6h+124\neq0\\
h\neq-\frac{124}{6}\\
h\neq-\frac{62}{3}\\"

If "h\neq-\frac{62}{3}, h\in(-\infty,-\frac{62}{3})\cup(-\frac{62}{3},\infty)" the vectors "\vec{a},\vec{b},\vec{c}" are linearly independent.